Part A: Carbonyl fluoride, COF2, is an important intermediate used in the production of fluorine-containing compounds. For instance, it is used to make the refrigerant carbon tetrafluoride, CF4 via the reaction 2COF2(g)?CO2(g)+CF4(g), Kc=5.10 If only COF2 is present initially at a concentration of 2.00 M, what concentration of COF2 remains at equilibrium? Express your answer with the appropriate units. Part B: Consider the reaction CO(g)+NH3(g)?HCONH2(g), Kc=0.770 If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?
Answer #1

Part a

The balanced reaction with ICE TABLE

2COF2(g)?CO2(g)+CF4(g)

I 2

C - 2x +x +x

E (2-2x) x x

Equilibrium constant expression of the reaction

Kc = [CF4] [CO2] /[COF2]2

5.10 = (x2) / (2-2x)2

2.258 = x/(2-2x)

2.258(2-2x) = x

4.516 - 4.516x = x

4.516 = 5.516x

x = 0.8187

At equilibrium

Concentration of COF2 = 2 - 2x = 2-2*0.8187 = 0.3626 M

Part b

CO(g)+NH3(g)?HCONH2(g)

I 1 2

C -x -x +x

E (1-x) (2-x) x

Equilibrium constant expression of the reaction

Kc = [HCONH2] /[CO] [NH3]

0.77 = x / (1-x)(2-x)

(0.77-0.77x)(2-x) - x = 0

1.54 - 0.77x - 1.54x + 0.77x2 - x = 0

0.77x2 - 3.31x + 1.54 = 0

x = 0.5308

concentration of HCONH2 at equilibrium = x = 0.5308 M